OpenStudy (anonymous):

Does the probability that you will draw a shape or a purple star involve disjoint or overlapping events? What formula would you use to calculate this probability?

OpenStudy (anonymous):

These are the cards I'm supposed to use.

OpenStudy (anonymous):

Please help! Will give medal and fan!

OpenStudy (phi):

A shape is any thing in the first row a purple star is the last element in the first row those events overlap.

OpenStudy (anonymous):

Could you explain more? I am still confused!?!

OpenStudy (phi):

You have 12 boxes. 4 of the boxes have shapes the probability of picking a shape is 4/12 or (simplifying) 1/4

OpenStudy (phi):

the purple star is a specific shape. if you picked a purple star you also picked a shape. That means the events: "pick a shape" and "pick a purple star" are not independent.

OpenStudy (anonymous):

Okay meaning that they are the same almost

OpenStudy (anonymous):

So all together what would my answer be?

OpenStudy (phi):

if you circle the events " a shape" you put a circle around the first row now circle the event " purple star" notice that the circles overlap. that means the events overlap

OpenStudy (phi):

to find the probability of overlapping events A or B use Pr( A or B) = Pr(A) + Pr(B) - Pr(A and B)

OpenStudy (phi):

The pr of a shape is 4/16 or 1/4

OpenStudy (phi):

the Prob of a purple star is 1/16 the prob of a purples star and a shape is still 1/16 so the answer is 1/4 + 1/16 - 1/16 = 1/4